### v1

# STANCE-$\text{C}^{3}$: Domain-adaptive Cross-target Stance Detection via Contrastive Learning and Counterfactual Generation

Abstract
Stance detection is the process of inferring a person’s position or standpoint on a specific issue to deduce prevailing perceptions toward topics of general or controversial interest, such as health policies during the COVID-19 pandemic. Existing models for stance detection are trained to perform well for a single domain (e.g., COVID-19) and a specific target topic (e.g., masking protocols), but are generally ineffectual in other domains or targets due to distributional shifts in the data. However, constructing high-performing, domain-specific stance detection models requires an extensive corpus of labeled data relevant to the targeted domain, yet such datasets are not readily available. This poses a challenge as the process of annotating data is costly and time-consuming. To address these challenges, we introduce a novel stance detection model coined domain-adaptive Cross-target STANCE detection via Contrastive learning and Counterfactual generation (STANCE-$\text{C}^{3}$) that uses counterfactual data augmentation to enhance domain-adaptive training by enriching the target domain dataset during the training process and requiring significantly less information from the new domain. We also propose a modified self-supervised contrastive learning as a component of STANCE-$\text{C}^{3}$ to prevent overfitting for the existing domain and target and enable cross-target stance detection. Through experiments on various datasets, we show that STANCE-$\text{C}^{3}$ shows performance improvement over existing state-of-the-art methods.
I Introduction
Inferring prevailing public attitudes toward specific personalities, policies, or events, especially those with controversial undertones, is one of the many applications of mining social media data. In natural language processing (NLP), the task of predicting the general public’s outlook for specific target keyword(s) is called stance detection. The target represents the subject or topic of interest, while the stance reflects the prevailing attitude or position. As an example from the tweet “Everyone should get the COVID-19 vaccine”, a positive stance is posited toward COVID-19 vaccine as the target. Stance detection has had diverse applications in tasks such as investigating users’ opinions on a new product [1], understanding public acceptance of a newly proposed legislation [2], and determining support for public health measures [3].
Stance detection models have been proposed under two constructs, namely, the single-target and cross-target scenarios. The single-target scenario determines users’ opinions toward a single, fixed target [4], while cross-target methods extend the model’s versatility to address multiple targets [5]. In single-target cases, the training set includes labeled data related to the target, while in cross-target methods, there is lacking or very few labeled data on the target of interest.
In addition, current state-of-the-art (SOTA) stance detection models are primarily focused on approaches for improving the model’s performance on a single domain. Previous research has illustrated, however, that these models do not necessarily perform well when used in a domain that is distinct from the domain of the training data set  [6]. We hypothesize that this performance degradation is mainly due to the following attributes of current SOTA stance detection models: (1) the tendency for attention bias toward events related to the given target, a phenomenon that we will illustrate in experiments in later sections; (2) the absence of target words in predicting stance during the training phase; and (3) the model’s reliance on auxiliary words (i.e., words that are not necessarily the target or stance) specific to a given domain. Additionally, the scarcity of available training data limits the capabilities of stance detection models across multiple domains and targets.
Table I shows an example of a domain-adaptive cross-target stance detection problem. In this example, the two domains differ by hashtags and data acquisition time periods, while the targets are distinguished by the topical content (vaccination vs. mask wearing). Other attributes of the sampling process could account for distinctions among domains and targets, such as data sources (e.g., social media vs. news) or content genre (political vs. gossip news).
TABLE I: Examples of domain-adaptive cross-target stance detection
Source Target:
COVID-19 vaccination
Statement: Got my first vaccine dose. I’m so thankful for all the healthcare workers that made the experience positive.
Stance: In favor of
Source Domain:
Tweets with target-related keywords (e.g. get vaccinated) from January 1st, 2020 to August 23rd, 2021
Destination Target:
wearing face masks
Statement: I will not take the mask and I will not wear a mask. In the words of Patrick Henry "Give me liberty or give me death!"
Stance: Against
Target Domain:
Tweets with target-related hashtags (e.g. #MasksSaveLives) from February 27th, 2020 to August 20th, 2020
In this paper, we propose a cross-domain, cross-target stance detection model called domain-adaptive Cross-target STANCE detection via Contrastive learning and Counterfactual generation (STANCE-$\text{C}^{3}$) to address limitations in existing SOTA models. STANCE-$\text{C}^{3}$ is a domain- and target-agnostic model that leverages counterfactual data augmentation (CDA) [7], an approach that has been shown to decrease bias and improve the robustness of neural network-based models [8]. We combine CDA with text style transfer methods [9] to facilitate effective text transfer from one domain to another. These language model-based data augmentation techniques are expected to enrich existing datasets with information that will improve the performance of stance detection models across multiple domains.
Previous work on stance detection addressed domain robustness by using adversarial training and including the target word as an input to the stance detection model (e.g., concatenating the target word and text as input to the language model). However, this approach resulted in overfitting the model to the specific target, thus limiting its performance robustness for other targets. In STANCE-$\text{C}^{3}$, we adopt a self-supervised technique called contrastive learning [10] to extract both shared and distinct high-level features that capture the specific characteristics of statement-target pairs. Text features such as word embeddings, Part-of-Speech tags, or $n$-grams are commonly used across tasks or datasets, while domain-specific terms, phrases, or concepts are normally used for target topics or subjects. In addition to the existing contrastive loss function, we incorporate additional components such as cosine similarity and negative pair loss to guide the model to minimize the distance between pairs with same stance labels and maximize the distance for pairs with different stance labels. Because contrastive learning is classified as a self-supervised approach, STANCE-$\text{C}^{3}$ benefits from requiring fewer manually annotated data points. Additionally, when combined with CDA during the pre-processing stage, it ensures that the training set is optimally constructed for a cross-domain and cross-target setting, i.e., there are sufficient data points exhibiting both similarities and differences, allowing the model to infer salient features that are useful for stance detection.
The major contributions of this work are as follows: (1) we investigate the problem of domain-adaptive cross-target stance detection and propose a novel solution for training a model when data on the target domain and target object (or destination target) are sparse; (2) we demonstrate the utility of contrastive learning as a possible solution for improving a model’s robustness across domains; and finally, (3) we provide empirical evidence of STANCE-$\text{C}^{3}$’s effectiveness by running experiments and ablation studies on real-world datasets.
II Problem Statement
Suppose we have two sets of instances $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{D}_{\mathcal{T}}$ from source domain $\mathcal{S}$ and target domain $\mathcal{T}$, respectively. The set $\mathcal{D}_{\mathcal{S}}={(r_{\mathcal{S}},t^{o}_{,}y_{\mathcal{S}})}^{N}$ contains N annotated instances from a source domain $\mathcal{S}$ where each instance $r_{\mathcal{S}}$ is a text (e.g., tweet) consisting of a sequence of k words, and $y_{\mathcal{S}}$ is the stance toward a target $t^{o}$ associated with each annotated instance. Another set $\mathcal{D}_{\mathcal{T}}=(r_{\mathcal{T}},t^{d})^{N^{\prime}}$ is a set of N’ instances from target domain $\mathcal{T}$ with the destination target $t^{d}$. The goal of the domain-adaptive cross-target stance detection is to learn the stance classifier $F$ that uses the features of the source domain $\mathcal{S}$ and target words $t^{o}$ in $\mathcal{D}_{\mathcal{S}}$ and predict the stance label $y_{\mathcal{T}}$ of each text $r_{\mathcal{T}}$ in $\mathcal{D}_{\mathcal{T}}$. We formally define the problem as follows:
Definition (Domain Adaptive Cross-Target Stance Detection). Given statements from two separate datasets $\mathcal{D}_{\mathcal{S}}$ and $\mathcal{D}_{\mathcal{T}}$, and corresponding targets $t^{o}$ and $t^{d}$ from the source $\mathcal{S}$ and target $\mathcal{T}$ domains, respectively, learn a domain and target-agnostic text representation using $\mathcal{D}_{\mathcal{S}}$ and a small portion of labeled $\mathcal{D}_{\mathcal{T}}$ that can be classified correctly by the stance classifier $F$.
III Proposed Model
Figure 1: The proposed STANCE-$\text{C}^{3}$ architecture consists of two main components: counterfactual data generation, and contrastive learning networks. First, the counterfactual data generation network aims to enhance generalization by augmenting the dataset. In this process, tokens are represented as numeric vectors and a T5-based language model is trained to generate examples with the same sentence structure but different semantic context (i.e. domain). These generated examples are referred to as domain counterfactuals. Then using the augmented data, the contrastive learning network learns a cross-target representation that can be used for domain-adaptive cross-target stance detection.
In this section, we describe our proposed model STANCE-$\text{C}^{3}$. As shown in Figure 1, the architecture consists of two main components: (1) a domain counterfactual generator that is trained with input representations of the labeled samples from the source and a small portion from the target domain and generates a set of domain counterfactuals. The generated counterfactuals provide knowledge of the target domain during the stance classification training phase. (2) These samples are then combined with the original dataset (i.e., source domain dataset with a small portion of target domain dataset) to train a BERT-based stance classifier using contrastive learning to create a domain-adaptive cross-target stance detection classifier. We split a small number of randomly selected samples from the combined datasets into positive and negative pairs according to their stance regardless of the target. This approach for contrastive loss enables us to improve the stance detection method’s generalization to unseen targets. In the following subsections, we explain the two stages in detail.
III-A Domain-adaptive Single-target
One major challenge in domain-adaptive stance detection is that collecting samples for a target domain can be expensive and usually requires human annotations. Even with annotated data, modeling a domain-adaptive stance detection is difficult due to the complexity of the dataset which contains far more distinct features over different domains [11]. These features often form spurious correlations within a specific domain and make the models brittle to domain shift. For example, the presence of question marks in a specific stance or highly negative sentiment associated with a particular stance leads to a decrease in performance [6].
To address the aforementioned challenges and control these variants on the domain-adaptive stance detection problem, we adapt counterfactual concepts to the source domain. A domain-counterfactual example is defined as a result text of intervening on the domain variable of the original example and converting it to the target domain while keeping other features (e.g., overall structure of the sentence) fixed. Given an example data from the COVID-19 vaccination domain (source domain), the generator first recognizes the terms related to the domain. Then, it intervenes on these terms, replacing them with text that links the example to the wearing face mask domain (target domain) while keeping the stance. We utilize source domain instances in combination with a small portion of target domain instances to incorporate target domain information in generating the counterfactuals specific to the target domain. These counterfactuals are then combined as input for the subsequent stage. Following the work of  [12], we adopt a T5-based language model (LM) to generate counterfactuals for a given target domain $\mathcal{T}$. In this approach, we train the T5-based LM to convert samples from source domain $\mathcal{S}$ to a given target domain $\mathcal{T}$.
Generating domain counterfactuals consists of two main steps - domain corruption and reconstruction. In the first step of domain corruption, we get the masking score of all n-grams $w$ ($n\in\{1,2,3\}$) in a dataset $\mathcal{D}_{\mathcal{S}}$ following [12]. The affinity of $w$ to domain $\mathcal{S}$ is defined as $\rho(w,\mathcal{S})=P(\mathcal{S}|w)(1-\frac{H(\mathcal{S}|w)}{\log N}),$ where $H(\mathcal{S}|w)$ is the entropy of $\mathcal{S}$ given each n-gram $w$ and $N$ is the number of unlabeled domains we know. The final masking score given source domain $\mathcal{S}$ and target domain $\mathcal{T}$ is $mask(w,\mathcal{S},\mathcal{T})=\rho(w,\mathcal{S})-\rho(w,\mathcal{T})$. Here, a higher score implies the n-gram is highly related to the source domain and distant from the target domain.
The next step is to reconstruct the masked examples $M(x)$ by concatenating them with a trainable embedding matrix, domain orientation vectors. The matrix is initialized with the embedding vector of each domain name and domain representative words. The concatenated matrix is trained with an encoder-decoder T5 architecture. Note that $\mathcal{T}=\mathcal{S}$ and $mask(w,\mathcal{S},\mathcal{T})=0$ during the training process. Given the target domain and its orientation vectors in the test phase, the trained model generates domain counterfactual $x^{\prime}$.
III-B Domain-adaptive Cross-target
One of the main challenges in detecting stances across different targets is that the distribution of data varies for each target even within the same domain. For instance, the word WHO may appear more frequently in the domain where the target is COVID-19 vaccination than in another domain where the target is Donald Trump [13]. The common occurrences between specific target words and instances make the model biased toward learning. Therefore it is necessary to identify an effective way to model transferable knowledge. To learn target-invariant features in instances from both the origin $t^{o}$ and destination $t^{d}$ targets, we use a simple yet effective supervised contrastive learning approach.
Contrastive learning has been widely used for many applications in Computer Vision, such as identity recognition tasks. The goal of contrastive learning is to pull the distance between an anchor and positive pair and push the distance between an anchor and negative pair. Here, we use the supervised contrastive loss [14], which applies a modified version of contrastive learning to supervised learning settings. The supervised contrastive loss function aims to enhance the separability of different classes by maximizing the distance between the anchor and the negative samples by pushing away the samples from different classes while minimizing the distance between the anchor and the positive samples by pulling together the data points from the same class.
In our problem setting, considering samples with different stance targets $t^{o}$ and $t^{d}$, the goal of the contrastive learning approach is to reduce the distance between samples with the same stance label and vice-versa. For example, we aim to maximize the distance between pairs of representations obtained from tweets that exhibit different stances towards the same target. Note that contrastive learning is applied to all pairs, regardless of whether the targets in the pair are equivalent or not. This approach creates a representation for the stance classifier that indicates the relation between the statement and the stance target $t$. We use the following modified supervised contrastive learning method in the spirit of [14]:
$\ell_{\mathrm{cont}}=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(z_{i}\cdot z_{p}/\tau)\cdot sim(z_{i},z_{p})}{\sum_{a\in A(i)}\exp(z_{i}\cdot z_{a}/\tau)}$
[…]
The loss takes a batch of samples $i\in I$ as an "anchor" from the training set where each sample is mapped to a vector $z_{i}$. Each anchor allows for multiple positive vectors $z_{p}$ that have the same label as an anchor ($y_{j}=y_{i}$) in addition to multiple negative vectors $z_{n}$ ($z_{n}\in A(i)$, $y_{n}\neq y_{i}$). $P(i)$ and $N(i)$ represent the indices of all positive and negative vectors, respectively. $\tau$ shows the temperature parameter that affects the distance of instance discrimination. The notation $sim(\cdot)$ refers to the cosine similarity. Here, we introduce the second term to balance the weight of positive and negative pairs. According to our observations on experiments involving different variants of the modified contrastive loss, we discovered that weighing the original loss with text similarity improves the model’s generalization ability across different targets. Note that anchor $z_{i}$ is from target $t^{o}$ but $z_{p}$ and $z_{n}$ are from both $t^{d}$ and $t^{o}$ for target-agnostic feature learning.
Optimization Algorithm The training process of STANCE-$\text{C}^{3}$ has two stages [^1] Source code and datasets will become publicly available upon acceptance.. During the first stage, the parameters of the T5-based LM are trained for conversion between source $\mathcal{S}$ and the target $\mathcal{T}$ domain following the work of Calderon et al. [12]. This stage enriches the input dataset to include more samples similar to the target domain. The next stage uses the contrastive learning approach to create a cross-target sentence representation. During the contrastive learning process, we add an additional loss $\ell_{\mathrm{cont}}$ to the classification loss for cross-target stance detection. For each instance $x_{i}$, the overall loss is formulated as $\ell_{\mathrm{total}}=\lambda\ell_{\mathrm{cont}}+(1-\lambda)\ell_{\mathrm{CE}}$ where $\ell_{\mathrm{CE}}$ stands for the cross-entropy loss for classification and $\lambda$ is the weighting factor. The cross-entropy loss is calculated as $CE=-\sum_{i=1}^{k}y_{i}\log(p_{i})$ for $k$ classes based on the true stance label $y$ and predicted stance probability $p$ for the $i^{th}$ class. Similar to [14] we set the temperature parameter $\tau=0.08$.
IV Experiments
The experiments conducted in this study investigate how changing the domains and targets affect the performance of several stance detection models, including STANCE-$\text{C}^{3}$. We also examine the effects of data augmentation and contrastive learning on performance accuracy in cross-domain and cross-target scenarios. More specifically, we aim to answer the following research questions: Q1. Does STANCE-$\text{C}^{3}$  achieve better performance in comparison to other SOTA when trained on source domain $\mathcal{S}$ with some data from target domain $\mathcal{T}$, and tested on target domain $\mathcal{T}$ for target words in the source domain $\mathcal{S}$? Q2. Does STANCE-$\text{C}^{3}$  achieve better performance in comparison to other SOTA when trained on a source domain $\mathcal{S}$ and tested on a target domain $\mathcal{T}$ for target words not in the source domain? Q3. What is the effect of changes in the model’s parameters and components on its performance in the stance detection task? We consider the following scenarios in the experimental design: (1) single domain + single target, (2) cross-domain + single target, and (3) cross-domain + cross-target, where the targets and domains of the datasets have clear distributional differences. We address Q1 by training and testing the baseline and proposed models under the first scenario where both domain and target are the same for training and testing sets. For Q2, the models are evaluated under scenarios (2) and (3) to compare and contrast the models’ domain-adaptive and cross-target performance. Finally, we address Q3 by performing ablation study and parameter analysis on datasets with different stance targets.
TABLE II: Baselines targeted goals indicate the difference between the proposed approach STANCE-$\text{C}^{3}$ and the baselines.
Model Cross-Domain Cross-Target
MoLE [15]
[…]
MTL [11]
[…]
REAL-FND [16]
[…]
MTSD [17]
[…]
IV-A Baselines
We consider the following baselines. Table II shows the difference between baselines and the proposed approach. Note that we have implemented the baselines for both cross-domain and cross-target tasks in the table, regardless of their specific task targets. This choice stems from the unique nature of the task, which presents limitations in terms of available baselines.
MoLE [15]: Utilizes mixture-of-experts models where each domain expert represents each domain and produces probabilities for all the target labels. Then uses domain adversarial training to learn domain-invariant representations.
MTL [11]: A benchmarking framework for evaluating the robustness of stance detection systems. The proposed method leverages a diverse set of challenging datasets with varying levels of noise, bias, and adversarial attacks to evaluate the performance and robustness of stance detection systems. For each dataset, this model adds a domain-specific projection layer to the final layer of a pre-trained language model and freezes other layers during training.
RLFND [16]: A domain-adaptive reinforcement learning framework for fake news detection. The proposed method leverages a deep reinforcement learning algorithm to learn the optimal policy for fake news detection while adapting to the domain shift between the source and target domains.
MTSD [17]: A multi-task learning framework that performs stance detection on multiple targets simultaneously. The framework leverages a shared encoder to capture the common features of the text and a task-specific decoder to predict the stance toward each target.
We compare our approach to the existing state-of-the-art approaches using accuracy and AUC metrics.
IV-B Datasets
TABLE III: Datasets’ statistics.
Dataset Target Source Labels Samples
CoVaxNet
COVID-19 Vaccination
Twitter
Pro, Anti
1,831,220
COVID-19-Stance
Face masks, Fauci, School closures, Stay at home orders
Twitter
Favor, Against, None
7,122
We use two representative datasets related to COVID-19 with different domains and targets (see Table III). The large-scale Twitter dataset CoVaxNet [18] was used for both cross-domain and cross-target training. CoVaxNet was acquired over the pandemic period (Jan 1st, 2020 - Dec 31st, 2021) and labeled with stances toward “COVID-19 vaccination”. For testing the performance of STANCE-$\text{C}^{3}$  in a single-target setting, the domains are defined as separate time periods marked by a changepoint event. We divided CoVaxNet into two time periods with the changepoint event occurring on August 23, 2021, when the US Food and Drug Administration (FDA) fully approved the Pfizer-BioNTech vaccine. Thus, the data for the source and target domains are 10,000 randomly selected tweets from Jan 1st, 2020 to August 22, 2021 (“CoVaxNet_pre”; source domain) and 10,000 tweets from August 23, 2021 to December 31st, 2021 (“CoVaxNet_post”; target domain). For the cross-target scenario, CoVaxNet_pre and CoVaxNet_post were separately used as training sets for STANCE-$\text{C}^{3}$. Model performance was tested on targets (“Anthony S. Fauci, M.D.”, “Wearing a Face Mask”, “Keeping Schools Closed”, “Stay at Home Orders”) from the COVID-19-Stance [19] dataset. COVID-19-Stance was collected in the period (February 27th, 2020 - August 20th, 2020) with different targets and keywords of interest (e.g., #lockdown), thus justifying an adequate distributional shift between the source and target. It should be noted that we treated the congruent labels from the two datasets as equivalent i.e., pro is equivalent to favor and anti is the same as against. In the context of stance detection, the words in these pairs essentially convey the same stance.
TABLE IV: Cross-domain, single-target performance results. In this experiment we only use the datasets that have the same target word (p-value < 0.05 for all McNemar’s tests).
Source Domain Target Domain BERT MoLE MTL MTSD RLFND STANCE-$\text{C}^{3}$
CoVaxNet_pre CoVaxNet_post 0.713 (0.697) 0.646 (0.621) 0.734 (0.736) 0.734 (0.731) 0.752 (0.747) 0.765 (0.753)
CoVaxNet_pre 0.762 (0.727) 0.901 (0.876) 0.823 (0.789) 0.845 (0.831) 0.843 (0.835) 0.863 (0.858)
Performance Degradation 0.049 (0.030) 0.255 (0.255) 0.089 (0.053) 0.111 (0.100) 0.091 (0.088) 0.098 (0.105)
CoVaxNet_post CoVaxNet_post 0.826 (0.796) 0.821 (0.835) 0.811 (0.823) 0.841 (0.839) 0.812 (0.784) 0.831 (0.837)
CoVaxNet_pre 0.698 (0.655) 0.713 (0.702) 0.623 (0.611) 0.712 (0.692) 0.674 (0.632) 0.723 (0.723)
Performance Degradation 0.128 (0.141) 0.108 (0.133) 0.188 (0.212) 0.129 (0.147) 0.138 (0.152) 0.108 (0.114)
TABLE V: Cross-domain, cross-target performance results. In this experiment, we evaluate the model’s performance on a dataset that has a different target word in comparison to the source domain’s target (p-value < 0.05 for all McNemar’s tests).
Source Domain Target Domain BERT MoLE MTL MTSD RLFND STANCE-$\text{C}^{3}$
CoVaxNet_pre Face Masks 0.623 (0.568) 0.732 (0.711) 0.741 (0.712) 0.789 (0.746) 0.742 (0.625) 0.868 (0.823)
Fauci 0.741 (0.706) 0.725 (0.719) 0.691 (0.687) 0.732 (0.698) 0.699 (0.625) 0.743 (0.714)
School Closures 0.832 (0.627) 0.778 (0.761) 0.711 (0.651) 0.872 (0.731) 0.834 (0.804) 0.920 (0.734)
Stay at Home 0.641 (0.625) 0.734 (0.738) 0.621 (0.613) 0.738 (0.732) 0.665 (0.658) 0.736 (0.742)
Average Performance 0.709 (0.631) 0.742 (0.732) 0.691 (0.665) 0.782 (0.726) 0.735 (0.678) 0.816 (0.753)
CoVaxNet_post Face Masks 0.570 (0.555) 0.801 (0.786) 0.698 (0.676) 0.794 (0.754) 0.734 (0.712) 0.819 (0.769)
Fauci 0.548 (0.624) 0.721 (0.740) 0.709 (0.683) 0.763 (0.778) 0.713 (0.694) 0.743 (0.714)
School Closures 0.578 (0.543) 0.718 (0.743) 0.621 (0.620) 0.687 (0.701) 0.694 (0.675) 0.739 (0.772)
Stay at Home 0.623 (0.637) 0.672 (0.691) 0.669 (0.681) 0.699 (0.704) 0.675 (0.662) 0.725 (0.727)
Average Performance 0.579 (0.589) 0.728 (0.740) 0.674 (0.665) 0.735 (0.734) 0.704 (0.685) 0.756 (0.745)
IV-C Evaluation and Results
Table IV and Table V summarize the comparative performance of the baseline and proposed models with respect to the accuracy and AUC (in parentheses).
Q1: In this experiment, the proposed approach and the domain-adaptive baselines used a fixed portion (30%) of the target domain data during counterfactual data generation. The models were trained on a single domain and tested on both source and target domains for stance detection for a single target of interest from the source domain. Our results provide some evidence that even when the target of interest is contained in the source domain, performance in the target domain generally degrades for all models. The Performance Degradation in the table highlights how a model’s performance changes when shifting from the source domain to the target domain. We observe that all of the models exhibit performance degradation which confirms the difference between domains.
Q2: The results of our experiments on the cross-domain cross-target setting, as presented in Table V, demonstrate the performance of our approach and the baselines. In this setting, we augment the training set of baselines with a small portion (30%) of data from the target domain. The results of our experiments suggest that the proposed approach is more effective than the domain-adaptive baselines. Specifically, our approach surpasses all baselines in terms of average accuracy and AUC score, underscoring its suitability for the cross-domain cross-target context. Furthermore, these results underscore the robustness of our approach to variations in the target domain, a strength attributable to our counterfactual data generation technique which is tailored to produce data that better represents the target domain.
Parameter Analysis To answer Q3, we analyze the effect of hyperparameters on the performance of STANCE-$\text{C}^{3}$. We show the effect of crucial training parameters $\lambda$ and the target domain portion $\gamma$ in the training set. To demonstrate the effect of contrastive loss and the balance between contrastive and cross-entropy loss, we vary the $\lambda$ value in the range of $\{0.0,0.25,0.5,0.75,1.0\}$. As illustrated in Figure 2-a, the AUC score varies for different $\lambda$ values. The results indicate that using $\lambda=0.5$ yields the best performance.
Next, we investigate the effect of the target domain data portion $\gamma$ used for counterfactual data generation. We vary the $\gamma$ value within the range of $\{5\%,15\%,30\%,45\%\}$. The results, as depicted in Figure 2-b, show that adding more target domain data leads to better performance. However, in real-world scenarios, access to a large amount of data in the target domain is often limited. Furthermore, there is no significant performance difference between $\gamma=30\%$ and $\gamma=45\%$. Therefore, we conclude that using $30\%$ of the target domain data enables STANCE-$\text{C}^{3}$ to achieve an acceptable performance.
[…]
V Related Work
STANCE-$\text{C}^{3}$ is proposed in this paper to improve existing models for stance detection and robustness across domains. In this section, we review current NLP SOTA models for stance detection and domain adaptation.
V-A Stance Detection
There are two general categories of distinctions among stance detection models. Firstly, models are distinguished on how stance features are modeled. Content-level approaches use linguistic features such as topics [20] or targets [21] alongside sentiment information [22]. User-level approaches focus on user-related attributes interactions [23], information [24], and timelines [25]. Hybrid models combine content- and user-level features for more comprehensive representations [26]. In this paper, we focus on content-level modeling with target embedding using BERT due to its practicality in data acquisition.
Secondly, stance detection models differ with respect to the specificity of the target of interest. While many previous works focus on specific single-target scenarios [27], recent studies have considered cross-target [21], multi-target [17], and few-shot or zero-shot stance detection [28]. While both cross-target and zero-shot stance detection models are trained on one or more source targets and used on previously unseen targets [29], zero-shot methods are able to predict stances for unseen targets that are not necessarily related to the target in the training set [30]. Both attempt to extract transferable knowledge from the source using methods such as hierarchical contrastive learning[31]. STANCE-$\text{C}^{3}$uses a similar but modified approach to the contrastive loss during training.
V-B Domain Adaptation
As a category of transfer learning, domain Adaptation (DA) leverages on agnostic or common information between domains with the target task remaining constant [32]. There are several DA setups, including unsupervised domain adaptation (UDA) where both labeled source data and unlabeled target data are available [33], semi-supervised domain adaptation (SSDA) where labeled source data and a small number of labeled target data are accessible [34], and supervised domain adaptation (SDA) where both labeled source and target data are available during training [35]. Our focus is on SSDA, where a smaller set of labeled target data is used in the training process.
Domain adaptive models are categorized by their focus on the model, training data, or both. Model-centric approaches aim to construct domain-agnostic model structures, such as in Deng et al. [36]. Data-centric methods enhance robustness by leveraging on data attributes using techniques such as the incorporation of loss functions based on inter-domain distances [37] or the employment of pre-training [38]. Finally, hybrid approaches combine model- and data-centric approaches [15]. Many of these DA methods grapple with spurious correlations from differing training and test dataset distributions, thus limiting their effectiveness. Our hybrid approach addresses this by using domain counterfactual examples, which minimizes the impact of domain-invariant features in cross-domain classification.
VI Conclusion
The robustness of NLP models across domains and multiple targets of interest remains an open area of work in stance detection tasks. In this paper, we proposed a stance detection model STANCE-$\text{C}^{3}$ that exhibited better performance over existing SOTA for cross-domain, multi-target scenarios. By conducting experiments using datasets from the COVID pandemic, we showed that cross-domain and cross-target performance generally degrades for all models. However, in comparison to models such as MoLE and MTL, the cross-domain degradation performance of STANCE-$\text{C}^{3}$ is significantly lower, providing some evidence of its robustness across domains. Additionally, STANCE-$\text{C}^{3}$ generally outperforms the SOTA for multiple, cross-domain targets. Ablation studies on STANCE-$\text{C}^{3}$’s novel components namely, the contrastive loss function and the use of domain counterfactuals for data augmentation, illustrate the marginal impacts of these components on performance and explain why STANCE-$\text{C}^{3}$performed better than its SOTA counterparts. Future work includes extending the proposed training strategy to parameter-efficient models on large language models, along with application to the zero-shot stance detection tasks and other benchmark datasets.
Acknowledgments
This material is based on work supported by the U. S. Department of Homeland Security under Grant Award Number 17STQAC00001-05-00222Disclaimer: ”The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Homeland Security.”.